# Rectified Diffusion Guidance for Conditional Generation

### v1

@@ 4 Experiments
4.1 Experimental Setups
Datasets and baselines. We apply ReCFG to previous seminal DPMs, including LDM (Rombach et al., 2022) on ImageNet 256 (Deng et al., 2009), EDM2 (Karras et al., 2024b) on ImageNet 512, and SD3 (Esser et al., 2024) on CC12M (Changpinyo et al., 2021), respectively.
Evaluation metrics. As for LDM and EDM2, we draw 50,000 samples for Fréchet Inception Distance (FID) (Heusel et al., 2017) to evaluate the fidelity of the synthesized images. We further use Improved Precision (Prec.) and Recall (Rec.) (Kynkäänniemi et al., 2019) to separately measure sample fidelity (Precision) and diversity (Recall). As for SD3, following the official implementation, we use CLIP Score (CLIP-S) (Radford et al., 2021; Hessel et al., 2021) and FID on CLIP features (Sauer et al., 2021) on 1,000 samples to evaluate conditional faithfulness and fidelity of the synthesized images, respectively. Both two metrics are evaluated on the MS-COCO validation split (Lin et al., 2015).
Implementation details. We implement ReCFG with NVIDIA A100 GPUs, and use pre-trained LDM [^2] https://github.com/CompVis/latent-diffusion, EDM2333https://github.com/NVlabs/edm2, and SD3444https://huggingface.co/stabilityai/stable-diffusion-3-medium-diffusers provided in official implementation. We reproduce all the experiments with official and more other configurations including NFEs and guidance strengths.
4.2 Results on Toy Example in Section 3.2
We first confirm the effectiveness of our method on toy data, as presented in Sec. 3.2. Given the closed-form expressions of score functions, we are able to precisely describe the distribution of both gamma-powered distribution $q_{0,\gamma}(\mathbf{x}_{0}|c)$ by native CFG and $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ by ReCFG. The theoretical and numerical DDIM-based simulation value of probability density functions of both $q_{0,\gamma}(\mathbf{x}_{0}|c)$ and $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ are shown in Fig. 2. It is noteworthy that native CFG drifts the expectation of $q_{0,\gamma}(\mathbf{x}_{0}|c)$ further away from the peak of the ground-truth $q_{0}(\mathbf{x}_{0}|c)$ as $\gamma$ becomes larger, consistent with Theorem 2. As a comparison, the peaks of $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ and $q_{0}(\mathbf{x}_{0}|c)$ coincide, while $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ is sharpened with smaller variance. Therefore, by adopting relaxation on coefficients $\gamma_{1}$ and $\gamma_{0}$ with specially proposed constraints, our ReCFG manages to annihilate expectation shift, enabling better guidance and thus better conditional fidelity.
Table 1: Sample quality on ImageNet (Deng et al., 2009). For clearer demonstration, settings of $\gamma_{1}+\gamma_{0}=1$ (i.e. falling into native CFG) are highlighted in gray.
@@ 4.1 Experimental Setups
$\gamma_{1}$ $\gamma_{0}$ NFE ()($\downarrow$)( ↓ ) FID ()($\downarrow$)( ↓ ) Prec. ()($\uparrow$)( ↑ ) Rec. ()($\uparrow$)( ↑ )
5.0 -4.0 20 18.87 0.95 0.15
5.0 ReCFG 20 16.95 0.91 0.18
3.0 -2.0 20 11.46 0.94 0.27
3.0 ReCFG 20 9.78 0.91 0.32
5.0 -4.0 10 16.78 0.94 0.16
5.0 ReCFG 10 14.46 0.89 0.22
3.0 -2.0 10 10.13 0.91 0.28
3.0 ReCFG 10 8.26 0.91 0.33
ImageNet 512x512, EDM2-S (Karras et al., 2024b)
$\gamma_{1}$ $\gamma_{0}$ NFE ()($\downarrow$)( ↓ ) FID ()($\downarrow$)( ↓ ) Prec. ()($\uparrow$)( ↑ ) Rec. ()($\uparrow$)( ↑ )
3.0 -2.0 63 6.81 0.85 0.43
3.0 ReCFG 63 5.59 0.84 0.43
2.5 -1.5 63 5.87 0.85 0.46
2.5 ReCFG 63 4.84 0.84 0.48
2.0 -1.0 63 4.18 0.85 0.52
2.0 ReCFG 63 3.61 0.84 0.52
1.4 -0.4 63 2.29 0.83 0.59
1.4 ReCFG 63 2.23 0.83 0.59
Table 2: Sample quality on CC12M (Changpinyo et al., 2021) For clearer demonstration, settings of $\gamma_{1}+\gamma_{0}=1$ (i.e. falling into native CFG) are highlighted in gray.
@@
$\gamma_{1}$ $\gamma_{0}$ NFE ()($\downarrow$)( ↓ ) CLIP-S ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ )
7.5 -6.5 25 0.268 72.24
7.5 ReCFG 25 0.270 71.83
5.0 -4.0 25 0.267 72.37
5.0 ReCFG 25 0.268 71.95
2.5 -1.5 25 0.262 70.50
2.5 ReCFG 25 0.263 69.99
7.5 -6.5 10 0.262 82.71
7.5 ReCFG 10 0.263 76.05
5.0 -4.0 10 0.268 72.55
5.0 ReCFG 10 0.269 70.31
2.5 -1.5 10 0.265 71.17
2.5 ReCFG 10 0.265 68.68
7.5 -6.5 5 0.209 156.60
7.5 ReCFG 5 0.229 140.89
5.0 -4.0 5 0.248 115.51
5.0 ReCFG 5 0.258 101.82
2.5 -1.5 5 0.261 101.95
2.5 ReCFG 5 0.263 96.80
4.3 Results on Real Datasets
We conduct extensive experiments on state-of-the-art DPMs to quantitatively convey the efficacy of ReCFG. Results are reported in Tabs. 2 and 2. We can tell that ReCFG is capable of better performance on both class-conditioned and text-conditioned DPMs under various guidance strengths and NFEs. Additionally, ReCFG achieves better CLIP-S especially on small NFEs and large guidance strength, indicating better conditional fidelity on open-vocabulary synthesis.
4.4 Analyses
Table 3: Variance of lookup table over condition $c$. Note that we employ pixel-wise lookup table involving timestep $t$. We report the the mean and variance of lookup table over $c$, which is computed by averaging on all timesteps $t$ and pixels.
@@ 4.4 Analyses
Expectation Ratio 1.0050 $\pm$ 0.0012 1.0060 $\pm$ 0.0119 1.0250 $\pm$ 0.0369 1.0125 $\pm$ 0.0281
Table 4: Ablation study of the number of traversals (the number after ReCFG) for lookup table on ImageNet (Deng et al., 2009). For clearer demonstration, baselines of native CFG are highlighted in gray.
@@
$\gamma_{1}$ $\gamma_{0}$ NFE ()($\downarrow$)( ↓ ) FID ()($\downarrow$)( ↓ ) Prec. ()($\uparrow$)( ↑ ) Rec. ()($\uparrow$)( ↑ )
3.0 -2.0 10 10.13 0.91 0.28
3.0 ReCFG-10 10 8.88 0.92 0.30
3.0 ReCFG-100 10 8.70 0.92 0.31
3.0 ReCFG-500 10 8.26 0.91 0.33
ImageNet 512x512, EDM2-S (Karras et al., 2024b)
$\gamma_{1}$ $\gamma_{0}$ NFE ()($\downarrow$)( ↓ ) FID ()($\downarrow$)( ↓ ) Prec. ()($\uparrow$)( ↑ ) Rec. ()($\uparrow$)( ↑ )
2.5 -1.5 63 5.87 0.85 0.46
2.5 ReCFG-10 63 5.06 0.84 0.47
2.5 ReCFG-100 63 4.99 0.84 0.45
2.5 ReCFG-500 63 4.84 0.84 0.48
2.0 -1.0 63 4.18 0.85 0.52
2.0 ReCFG-10 63 3.70 0.84 0.52
2.0 ReCFG-100 63 3.66 0.84 0.52
2.0 ReCFG-500 63 3.61 0.84 0.52
Table 5: Ablation study of the number of traversals (the number after ReCFG) for lookup table on CC12M (Changpinyo et al., 2021). For clearer demonstration, baselines of native CFG are highlighted in gray.
@@
$\gamma_{1}$ $\gamma_{0}$ NFE ()($\downarrow$)( ↓ ) CLIP-S ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ )
5.0 -4.0 25 0.267 72.37
5.0 ReCFG-10 25 0.267 72.15
5.0 ReCFG-100 25 0.268 72.03
5.0 ReCFG-500 25 0.268 71.95
5.0 -4.0 10 0.268 72.55
5.0 ReCFG-10 10 0.268 71.61
5.0 ReCFG-100 10 0.268 70.64
5.0 ReCFG-500 10 0.269 70.31
5.0 -4.0 5 0.248 115.51
5.0 ReCFG-10 5 0.252 107.09
5.0 ReCFG-100 5 0.256 103.25
5.0 ReCFG-500 5 0.258 101.82
Variance of lookup table over condition $c$. Note that we need to pre-compute the lookup table consisting of expectation ratios for all conditions $c$, which is time-consuming and impractical for open-vocabulary distributions (e.g., text-conditioned DPMs). In Tab. 3 we report the mean and variance of expectation ratios over condition $c$, which is averaged on all timesteps and pixels. It is noteworthy that the variance of text-conditioned DPMs is larger than that of class-conditioned ones, while both of which is insignificant compared to the mean. Therefore, it is feasible to prepare the lookup table for only part of all potential conditions and use the mean for all conditions, serving as a practical strategy to improve time efficiency.
Ablation studies. Recall that we pre-compute the lookup table by traversing the dataset $q_{0}(\mathbf{x}_{0}|c)$. We conduct comprehensive ablation studies to convey a direct and clear picture of the efficacy of ReCFG under different numbers of traversals, as reported in Tabs. 5 and 5. We can conclude that larger number of traversals suggests better guidance performance, yet improvements from 100 to 500 traversals are relatively inconspicuous. In other words, employment of 500 samples per condition is adequate to serve as an empirical setting.
Pixel-wise lookup table. ReCFG enables pixel-specific guidance coefficients $\gamma_{1}$ and $\gamma_{0}$, thanks to the closed-form solution to Eq. 33, i.e., we can assign $\gamma_{0}$ for each pixel by maintaining the lookup table of pixel-wise expectation ratios. Fig. 1 demonstrates the ratios on LDM, EDM2, and SD3 under different NFEs. It is noteworthy that there appears no general rules on the relation between $\gamma_{1}$ and $\gamma_{0}$, indicating that trivially setting $\gamma_{1}$ and $\gamma_{0}$ to be scalars is less reasonable. As a comparison, our method makes it possible to employ more precise control on guided sampling in a simple and post-hoc fashion without further fine-tuning, enabling better performance.
@@ 4.5 Discussions
Classifier-Free Guidance is designed from Bayesian theory to facilitate conditional sampling, yet appears incompatible with original diffusion theory. Therefore, we believe ReCFG is attached to great importance on guided sampling by fixing the theoretical flaw of CFG. Despite the success on better conditional fidelity, our algorithm has several potential limitations. We need to pre-compute the lookup table by traversing the dataset to achieve rectified coefficients for each condition. Although we conduct extensive ablation studies on the number of traversals and variance over condition $c$, providing an adequate strategy especially for open-vocabulary datasets on text-conditioned synthesis, the optimal strategy is unexplored. Besides, we at present cannot provide precise control on variance of ReCFG, and turn to employ empirical values. Therefore, how to further conquer these problems (e.g., employing a predictor network $\bm{\omega}(c,t)$ for better $\gamma_{0}$ on open-vocabulary datasets according to Eq. 37) will be an interesting avenue for future research. Although leaving the variance behavior unexplored, we hope that ReCFG will encourage the community to close the gap in the future.
@@ A.1 Proof of Theorem 1
[…]
@@
Similar to derivation in DDIM (Song et al., 2021), first rewrite $J_{\delta,\gamma}$ as below:
$J_{\delta,\gamma}=\mathbb{E}\left[-\log p_{\theta,\gamma}(\mathbf{x}_{0}|\mathbf{x}_{1},c)+\sum_{t=2}^{T}D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|p_{\theta,\gamma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))\right]+C_{1},$ (S19)
@@
$\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|p_{\theta,\gamma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))]$ (S20)
$=$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta,\gamma}^{t}(\mathbf{x}_{t},c),c))]$ (S21)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\mathbf{f}_{\theta,\gamma}^{t}(\mathbf{x}_{t},c)\|_{2}^{2}]$ (S22)
@@
$\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[-\log p_{\theta,\gamma}(\mathbf{x}_{0}|\mathbf{x}_{1},c))]$ (S29)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\mathbf{f}_{\theta,\gamma}^{t}(\mathbf{x}_{1},c)\|_{2}^{2}]+C_{2}$ (S30)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=f(t)\mathbf{x}_{t}-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c),$ (S33)
@@
$\mathbb{E}_{\mathbf{x}_{t}}[\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma}(\mathbf{x}_{t}|c)]$ $=\mathbb{E}_{\mathbf{x}_{t}}[\gamma\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})]$ (S37)
@@
Note that $\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\mathbf{x}_{t}]=\alpha_{t}\mathbb{E}_{q_{0}(\mathbf{x}_{0}|c)}[\mathbf{x}_{0}]$, and that $\mathbb{E}_{\mathbf{x}_{0},c}[\mathbf{x}_{0}]=\int\mathbb{E}_{q_{0}(\mathbf{x}_{0}|c)}[\mathbf{x}_{0}]\mathrm{d}c$. Therefore when $\gamma\neq 1$, $\mathbb{E}_{\mathbf{x}_{t}}[\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma}(\mathbf{x}_{t}|c)]$ is not guaranteed to be identical with 0. In other words, denoising with CFG cannot be expressed as a reciprocal of diffusion process with Gaussian noise. ∎

### v2

@@ 4 Experiments
Table 1: Sample quality on ImageNet [6].
@@
Table 2: Sample quality on CC12M [2]
@@
Table 3: Variance of lookup table over condition $c$ and timestep $t$. Note that we employ pixel-wise lookup table involving both $c$ and $t$. We report the the mean and variance of lookup table over $c$ and $t$, respectively, which are computed by averaging on all pixels.
@@
4.1 Experimental Setups
Datasets and baselines. We apply ReCFG to seminal class-conditioned and text-conditioned DPMs, including LDM [25] and DiT [21] on ImageNet 256 [6], EDM2 [16] on ImageNet 512, and SD3 [9] on CC12M [2], respectively.
[…]
Implementation details. We implement ReCFG with NVIDIA A100 GPUs, and employ pre-trained LDM [^2] https://github.com/CompVis/latent-diffusion, DiT [^3] https://github.com/facebookresearch/DiT, EDM2444https://github.com/NVlabs/edm2, and SD3555https://huggingface.co/stabilityai/stable-diffusion-3-medium-diffusers checkpoints provided in official implementation. We reproduce all the experiments with official and more other configurations including NFEs and guidance strengths.
4.2 Results on Toy Example in Section 3.2
We first confirm the effectiveness of our method on toy data, as presented in Sec. 3.2. Given the closed-form expressions of score functions, we are able to precisely describe the distributions of both gamma-powered distribution $q_{0,\gamma}(\mathbf{x}_{0}|c)$ by native CFG and $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ by our ReCFG. The theoretical and numerical DDIM-based simulation value of probability density functions of both $q_{0,\gamma}(\mathbf{x}_{0}|c)$ and $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ are shown in Fig. 1. It is noteworthy that native CFG drifts the expectation of $q_{0,\gamma}(\mathbf{x}_{0}|c)$ further away from the peak of the ground-truth $q_{0}(\mathbf{x}_{0}|c)$ as $\gamma$ becomes larger, consistent with Theorem 2. As a comparison, the peaks of $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ and $q_{0}(\mathbf{x}_{0}|c)$ coincide, while $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ is sharpened with smaller variance. Therefore, by adopting relaxation on coefficients $\gamma_{1}$ and $\gamma_{0}$ with specially proposed constraints, our ReCFG manages to annihilate expectation shift, enabling better guidance and thus better conditional fidelity.
4.3 Results on Real Datasets
We showcase some results in Fig. 3. One can see that ReCFG could fix artifacts on EDM2. It is also noteworthy that ReCFG significantly improves synthesis quality on SD3, especially detailed textures. Beyond the exhibited visualization, we conduct extensive quantitative experiments on state-of-the-art DPMs to further convey the efficacy of ReCFG. From Tabs. 2 and 2, we can tell that ReCFG is capable of better performance on both class-conditioned and text-conditioned DPMs under various guidance strengths and NFEs especially CLIP-S, indicating better conditional fidelity on open-vocabulary synthesis. Furthermore, correction for theoretical flaws of CFG enables strong compatibility of ReCFG with other empirical strategies such as RescaleCFG [19], achieving better performance.
4.4 Analyses
Table 4: Ablation study of the number of traversals (the number after ReCFG) for lookup table on ImageNet [6]. For clearer demonstration, baselines of native CFG are highlighted in gray.
@@ 4.4 Analyses
Table 5: Ablation study of the number of traversals (the number after ReCFG) for lookup table on CC12M [2]. For clearer demonstration, baselines of native CFG are highlighted in gray.
@@
Variance of lookup table. Note that we need to pre-compute the lookup table consisting of expectation ratios for all conditions $c$, which is time-consuming and impractical for open-vocabulary distributions (e.g., text-conditioned DPMs). In Tab. 3 we report the mean and variance of expectation ratios over condition $c$, which is averaged on all timesteps and pixels. One can observe that larger NFE suggests smaller ratio with also smaller variance. It is also noteworthy that the variance of text-conditioned DPMs is larger than that of class-conditioned ones due to far more complex open-vocabulary conditions, while both of which is insignificant compared to the mean. Therefore, it is feasible to prepare the lookup table for only part of all potential conditions and use the mean for all conditions, serving as a practically adequate strategy to improve time efficiency. Variance over timestep $t$ averaged on all pixels and part of conditions is also reported in Tab. 3, where similar conclusion could be achieved.
Ablation studies. Recall that we pre-compute the lookup table by traversing $q_{0}(\mathbf{x}_{0}|c)$ for each condition $c$. Comprehensive ablation studies reported in Tabs. 5 and 5 convey a direct and clear picture of the efficacy of ReCFG under different numbers of traversals. We can conclude that larger number of traversals suggests better guidance performance, yet improvements from 100 to 500 traversals are relatively inconspicuous, especially on text-conditoned DPMs. In other words, employing 500 samples per condition is adequate in practice to serve as an empirical setting.
Time cost. Given the analyses on variance of expectation ratios over condition $c$ and ablation on traversals, preparing the lookup table is quite efficient. In practice, we sample 500 images for a subset of 100 conditions, which takes $\sim 3$ hours using 1 NVIDIA A100 GPU. The time cost is very close to performing FID evaluation using 50,000 images.
Pixel-wise lookup table. ReCFG enables pixel-specific guidance coefficients $\gamma_{1}$ and $\gamma_{0}$ with the same shape as score functions, thanks to the closed-form solution in Eq. 42, i.e., we can assign $\gamma_{0}$ for each pixel by maintaining the lookup table of pixel-wise expectation ratios. Fig. 2 demonstrates the ratios on LDM, EDM2, and SD3 at uniformly sampled timesteps under different NFEs. Both LDM and EDM2 show the generation of class 0 in ImageNet (i.e., “tench”), while SD3 adopts the prompt “A bicycle replica with a clock as the front wheel”. One can observe that expectation ratios at the same timestep vary largely by different pixels, and there appears no general rules on the relation between $\gamma_{1}$ and $\gamma_{0}$. Therefore, it is indicated that trivially setting $\gamma_{1}$ and $\gamma_{0}$ to be scalars is less reasonable. As a comparison, our method makes it possible to employ more precise control on guided sampling in a simple and post-hoc fashion without further fine-tuning, enabling better performance. It is also noteworthy that expectation ratio of SD3 exhibits noticeable shapes, probably due to more informative text prompts than one-hot class labels and more powerful model thanks to the training scale.
@@ 4.5 Discussions
Classifier-Free Guidance is designed from Bayesian theory to facilitate conditional sampling, yet appears incompatible with original diffusion theory. Therefore, we believe ReCFG is attached to great importance on guided sampling by fixing the theoretical flaw of CFG. Despite the success on better conditional fidelity, our algorithm has several potential limitations. Theoretically, we need to pre-compute the lookup table by traversing the dataset to achieve rectified coefficients for each condition. Although we conduct extensive ablation studies on the number of traversals and variance over condition $c$, providing an adequate strategy especially for open-vocabulary datasets on text-conditioned synthesis, the optimal strategy is unexplored. Besides, we at present cannot provide precise control on variance of ReCFG due to incomputable variance in denoising process, and turn to employ empirical values. Therefore, how to further conquer these problems (e.g., employing a predictor network $\bm{\omega}(c,t)$ for better $\gamma_{0}$ on open-vocabulary datasets according to Eq. 45) will be an interesting avenue for future research. Although leaving the variance behavior unexplored, we hope that ReCFG will encourage the community to close the gap in the future.
@@ 5 Conclusion
Acknowledgements
This work was partially supported by the Natural Science Foundation of China (62461160309, 62302297), Beijing Science and Technology plan project (Z231100005923029), NSFC-RGC Joint Research Scheme (N_HKU705/24), Ant Group Research Intern Program, and Shanghai Sailing Program (22YF1420300).
Supplementary Material
Mengfei Xia1,2111Work finished during internship at Ant Group.   Nan Xue2   Yujun Shen2222Corresponding author.   Ran Yi3   Tieliang Gong4   Yong-Jin Liu1222Corresponding author.
1Tsinghua University   2Ant Group
3Shanghai Jiao Tong University   4Xi’an Jiao Tong University
@@ A.1 Proof of Theorem 1
in which $\langle\cdot,\cdot\rangle$ is inner product.
@@
Similar to derivation in DDIM [28], first rewrite $J_{\delta,\gamma}$ as below:
$J_{\delta,\gamma}=\mathbb{E}\left[-\log\hat{p}_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1},c)+\sum_{t=2}^{T}D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|\hat{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))\right]+C_{1},$ (S19)
@@
$\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|\hat{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c))]$ (S20)
$=$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[D_{KL}(q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)\|q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c),c))]$ (S21)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{t},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c)\|_{2}^{2}]$ (S22)
@@
$\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[-\log\hat{p}_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1},c))]$ (S29)
$\propto$ $\;\mathbb{E}_{q(\mathbf{x}_{1},\mathbf{x}_{0}|c)}[\|\mathbf{x}_{0}-\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{1},c)\|_{2}^{2}]+C_{2}$ (S30)
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$ $=f_{t}\mathbf{x}_{t}-\frac{1}{2}g^{2}_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c),$ (S33)
@@
$\mathbb{E}_{\mathbf{x}_{t}}[s_{t,\gamma}(\mathbf{x}_{t},c)]$ $=\mathbb{E}_{\mathbf{x}_{t}}[\gamma\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})]$ (S37)
@@
Note that $\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\mathbf{x}_{t}]=\alpha_{t}\mathbb{E}_{q_{0}(\mathbf{x}_{0}|c)}[\mathbf{x}_{0}]$, and that $\mathbb{E}_{\mathbf{x}_{0},c}[\mathbf{x}_{0}]=\int\mathbb{E}_{q_{0}(\mathbf{x}_{0}|c)}[\mathbf{x}_{0}]\mathrm{d}c$. Therefore when $\gamma\neq 1$, $\mathbb{E}_{\mathbf{x}_{t}}[s_{t,\gamma}(\mathbf{x}_{t},c)]$ is not guaranteed to be identical with 0. In other words, denoising with CFG cannot be expressed as a reciprocal of diffusion process with Gaussian noise. ∎